# Massive-STEPS: Massive Semantic Trajectories for Understanding POI Check-ins – Dataset and Benchmarks

### v2

@@ Appendix B Data Visualization
In Fig. 3, we show the top 10 most frequent POI categories for each city. The distribution reflects the local culture and lifestyle across different urban areas. For example, Beijing and Shanghai have a high number of Chinese restaurants, while Melbourne and Sydney show a strong presence of cafes. In Tokyo, convenience stores and ramen shops dominate. These patterns illustrate the diversity of local culture and user interests.
Fig. 4 plots the distribution of trajectory lengths (i.e., number of check-ins per trajectory). The distribution is long-tailed, with most trajectories being relatively short, similar to the original Semantic Trails dataset. This indicates that users often make only a few check-ins per outing.
Finally, we show the distribution of user activity levels, measured by the number of trajectories per user in Fig. 5. Most users exhibit cold-start behavior, contributing only a small number of trajectories. This highlights the importance of models that are robust to sparse and short user histories.
Table 6: Fields available in the Massive-STEPS dataset, including user, POI, geographic/spatial, and temporal details, along with example data for each field.
@@
Figure 3: Top 10 most frequent POI categories in each city, highlighting local cultural and urban preferences.
Figure 4: Distribution of trail lengths, showing a long-tailed pattern with most trajectories consisting of a few check-ins.
Figure 5: Distribution of user activity based on the number of trajectories per user, indicating a cold-start-heavy dataset.
@@ Appendix C POI Recommendation: Task Details
We adopt the conventional problem formulation used in prior POI recommendation studies  [68, 60, 54], which defines the task as learning user preferences and routines from historical check-ins to recommend future POIs.
C.1 Problem Formulation
@@
C.2 Experiment and Implementation Details
For training and evaluation, we used the LibCity [^1] https://github.com/libcity/bigscity-libcity-datasets/ library [41], which provides implementations of classical baselines including FPMC [32], RNN [41], LSTPM [35], and DeepMove [8]. The training hyperparameters are listed in Table 7 and, unless otherwise noted, follow the default configurations provided by LibCity.
For GETNext [^2] https://github.com/songyangme/GETNext [60] and STHGCN [^3] https://github.com/alipay/Spatio-Temporal-Hypergraph-Model [54], we adapted the original source code released by the respective authors. Due to variations in dataset sizes and training costs across cities, we applied different hyperparameters for some cities, as detailed in Table 8.
Table 7: Hyperparameters for Markov-based methods and recurrent networks baselines.
Hyperparameter FPMC RNN LSTPM DeepMove
Batch Size 20 20 20 20
Learning Rate 5e-4 1e-3 1e-4 1e-3
Max Epoch 1 30 40 30
Location Embedding Size 64 500 500 500
Hidden Embedding Size N/A 500 500 500
Dropout N/A 0.3 0.8 0.5
Table 8: Hyperparameters for Transformer-based graph neural networks.
Model Cities Batch Size LR Epochs
GETNext Beijing, Melbourne, Moscow, Shanghai, Sydney, Tokyo 16 1e-3 200
Istanbul, Kuwait City, New York, Petaling Jaya, São Paulo 16 1e-4 20
@@ C.2 Experiment and Implementation Details
STHGCN Beijing, Melbourne, Shanghai, Sydney, Tokyo 16 1e-4 20
Istanbul, Jakarta, Kuwait City, Moscow, New York, Petaling Jaya, São Paulo 64 1e-4 20
@@
All modified code implementations are available as submodules in our main dataset repository. Experiments were conducted using GPUs provided by the School of Computer Science and Engineering at UNSW Sydney via the Wolfpack computational cluster, which includes NVIDIA L4, L40S, and H100 GPUs.
C.3 Supplementary Results
@@
We report the full results of our supervised POI recommendation baselines in Table 9 and 10, using three evaluation metrics: Acc@1, Acc@5, and NDCG@5.
@@
Table 9: Performance of supervised POI recommendation baselines across 6 cities: Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, and Moscow. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model Beijing Istanbul Jakarta Kuwait City Melbourne Moscow
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.000 0.021 0.009 0.026 0.074 0.050 0.029 0.085 0.058 0.021 0.089 0.054 0.062 0.147 0.107 0.059 0.129 0.094
RNN 0.085 0.183 0.134 0.077 0.178 0.130 0.049 0.115 0.083 0.087 0.203 0.146 0.059 0.105 0.083 0.075 0.164 0.122
LSTPM 0.127 0.211 0.169 0.142 0.286 0.217 0.099 0.210 0.157 0.180 0.362 0.275 0.091 0.204 0.150 0.151 0.300 0.229
DeepMove 0.106 0.261 0.190 0.150 0.298 0.228 0.103 0.212 0.160 0.179 0.360 0.274 0.083 0.179 0.134 0.143 0.283 0.217
GETNext 0.433 0.527 0.486 0.146 0.268 0.210 0.155 0.257 0.209 0.175 0.322 0.251 0.100 0.250 0.179 0.175 0.335 0.260
STHGCN 0.453 0.640 0.552 0.241 0.385 0.318 0.197 0.334 0.270 0.225 0.394 0.314 0.168 0.318 0.247 0.223 0.382 0.308
@@
Table 10: Performance of supervised POI recommendation baselines across 6 cities: New York, Petaling Jaya, São Paulo, Shanghai, Sydney, and Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model New York Petaling Jaya São Paulo Shanghai Sydney Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.032 0.090 0.061 0.026 0.084 0.057 0.030 0.079 0.055 0.084 0.154 0.120 0.075 0.180 0.131 0.176 0.291 0.239
RNN 0.061 0.119 0.092 0.064 0.148 0.107 0.097 0.191 0.147 0.055 0.120 0.090 0.080 0.164 0.125 0.133 0.254 0.197
LSTPM 0.099 0.206 0.155 0.099 0.222 0.163 0.158 0.319 0.243 0.099 0.195 0.149 0.141 0.265 0.206 0.225 0.394 0.315
DeepMove 0.097 0.195 0.149 0.112 0.234 0.175 0.160 0.310 0.240 0.085 0.168 0.128 0.129 0.240 0.188 0.201 0.362 0.288
GETNext 0.134 0.263 0.202 0.139 0.254 0.200 0.202 0.360 0.286 0.115 0.230 0.177 0.181 0.347 0.266 0.180 0.361 0.275
STHGCN 0.146 0.259 0.207 0.174 0.301 0.241 0.250 0.425 0.344 0.193 0.329 0.264 0.227 0.378 0.307 0.250 0.432 0.350
@@ Appendix D Analyzing Urban Features and POI Recommendation Performance
As discussed in Section 2.2, several hypotheses have been proposed to explain why POI recommendation models perform better in certain cities than others. These hypotheses aim to uncover how various urban features affect model performance. For example, Gowalla-CA [4, 62] often yields lower accuracy compared to FSQ-NYC and FSQ-TKY [57], suggesting that some cities may be inherently harder to model. In this analysis, we focus on supervised models only.
Prior studies [60, 54, 22] have suggested several features as potential explanatory variables, including:
@@
Moreover, previous studies have primarily focused on only three datasets: FSQ-NYC, FSQ-TKY, and Gowalla-CA. In contrast, Massive-STEPS provides broader coverage across 12 cities, enabling a more comprehensive and robust analysis. To examine the relationship between urban features and model performance, we averaged the three evaluation metrics across six supervised baselines for each city and computed the Spearman correlation with each candidate feature. To further support our findings, we also included the results of GETNext [60] and STHGCN [54] on FSQ-NYC, FSQ-TKY, and Gowalla-CA, calculated their corresponding urban features, and averaged the reported metrics of each city. Fig. 6 presents the correlations between all nine features and the average performance metric.
Among all features, category entropy shows the strongest correlation with model performance, with a Spearman correlation of $r=-0.736$ ($p=0.002$). This suggests that cities with more evenly distributed POI categories tend to yield lower prediction accuracy. Intuitively, when no single category dominates (a city has roughly equal proportions of restaurants, cafes, homes, and other POIs), it becomes more difficult for models to predict a user’s next destination. In these cases, user behavior is more varied and less predictable. On the other hand, cities with more skewed category distributions (e.g., mostly food places or mostly residential areas) tend to have more consistent patterns of movement, making them easier for models to learn and predict. Interestingly, our finding contradicts the hypothesis proposed by LLM4POI [22], which suggests that FSQ-NYC is easier to model than Gowalla-CA due to the former’s vast number of POI categories, which were supposed to provide richer contextual signals for the model. Our results indicate that it is not the number of categories that matters, but rather how these categories are distributed.
Figure 6: Spearman correlation between nine candidate urban features and the mean score of POI recommendation models across 15 cities, including Massive-STEPS (ours), FSQ [57], and Gowalla [4, 62].
@@ E.1 Problem Formulation
E.2 Experiment and Implementation Details

### v3

@@ Appendix B Data Visualization
In Fig. 9, we show the top 10 most frequent POI categories for each city. The distribution reflects the local culture and lifestyle across different urban areas. For example, Beijing and Shanghai have a high number of Chinese restaurants, while Melbourne and Sydney show a strong presence of cafés. In Tokyo, convenience stores and ramen shops dominate. These patterns illustrate the diversity of local culture and user interests. Fig. 10 illustrates the temporal shift in the distribution of the top 10 most visited POI categories across each city, comparing the percentage of visits to each category between the 2012-2013 and 2017-2018 periods.
Figure 8. World map highlighting the cities included in the Massive-STEPS dataset.
World map highlighting the cities included in the Massive-STEPS dataset
Table 9. Fields available in the Massive-STEPS dataset, including user, POI, geographic/spatial, and temporal details, along with example data for each field.
@@
Figure 9. Top 10 most frequent POI categories in each city, highlighting local cultural and urban preferences.
Top 10 most frequent POI categories in each city, highlighting local cultural and urban preferences.
Figure 10. Top 10 most visited POI categories in each city across two time periods, illustrating temporal shifts in user visitation patterns.
Top 10 most visited POI categories in each city across two time periods, illustrating temporal shifts in user visitation patterns.
@@ Appendix C POI Recommendation: Task Details
We adopt the conventional problem formulation used in prior POI recommendation studies (Zhang et al., 2025; Yang et al., 2022c; Yan et al., 2023), which defines the task as learning user preferences and routines from historical check-ins to recommend future POIs.
C.1. Problem Formulation
@@
C.2. Models
For thoroughness, we evaluated the following models as baselines:
  • FPMC (Rendle et al., 2010): A classical baseline that combines first-order Markov chains with matrix factorization to model personalized next-location predictions.
  • RNN (Wang et al., 2021a), LSTPM (Sun et al., 2020), DeepMove (Feng et al., 2018), Flashback (Yang et al., 2021), Graph-Flashback (Rao et al., 2022), and LoTNext (Xu et al., 2024): Recurrent neural networks designed to capture sequential dependencies, with varying mechanisms to incorporate spatio-temporal context.
  • GETNext (Yang et al., 2022c) and STHGCN (Yan et al., 2023): Transformer-based graph neural networks to model social, spatial, and temporal dependencies.
  • UniMove (Han et al., 2025b): Trajectory foundation model based on a Transformer decoder architecture with Mixture of Experts (MoE) layers.
C.3. Experiment and Implementation Details
For training and evaluation, we used the LibCity [^4] https://github.com/libcity/bigscity-libcity-datasets/ library (Wang et al., 2021a), which provides implementations of classical baselines including FPMC (Rendle et al., 2010), RNN (Wang et al., 2021a), LSTPM (Sun et al., 2020), and DeepMove (Feng et al., 2018). The training hyperparameters are listed in Table 10 and, unless otherwise noted, follow the default configurations provided by LibCity.
For Flashback [^5] https://github.com/exascaleinfolab/flashback_code/ (Yang et al., 2021), Graph-Flashback [^6] https://github.com/kevin-xuan/Graph-Flashback (Rao et al., 2022), LoTNext [^7] https://github.com/Yukayo/LoTNext (Xu et al., 2024), GETNext [^8] https://github.com/songyangme/GETNext (Yang et al., 2022c), and STHGCN [^9] https://github.com/alipay/Spatio-Temporal-Hypergraph-Model (Yan et al., 2023), we adapted the source code released by the respective authors. Due to variations in dataset sizes and training costs across cities, we applied different hyperparameters for some cities, as detailed in Table 11.
For UniMove [^10] https://github.com/tsinghua-fib-lab/unimove (Han et al., 2025b), we modified their original source code for Massive-STEPS. For location features, we used Schema.org’s 162 list of categories as a categorical feature and the administrative region as the grid area for POI category distribution. Hyperparameters are listed in Table 10 and, unless otherwise noted, follow the default values.
Table 10. Hyperparameters for Markov-based methods, recurrent networks, and UniMove.
Hyperparameter FPMC RNN LSTPM DeepMove UniMove
Batch Size 20 20 20 20 4
Learning Rate 5e-4 1e-3 1e-4 1e-3 3e-4
Max Epoch 1 30 40 30 50
Location Embedding Size 64 500 500 500 {256, 128}
Hidden Embedding Size N/A 500 500 500 512
Dropout N/A 0.3 0.8 0.5 N/A
Table 11. Hyperparameters for Transformer-based graph neural networks.
Model Cities BS LR Epochs
Flashback, Beijing 10 1e-2 100
Graph-Flashback, Bandung, Istanbul, Jakarta, KC, Melbourne, Moscow, NY 100 1e-2 100
LoTNext Palembang, PJ, SP, Shanghai, Sydney, Tangerang, Tokyo 100 1e-2 100
GETNext Beijing, Melbourne, Moscow, Palembang, Shanghai, Sydney, Tokyo 16 1e-3 200
Bandung, Istanbul, KC, NY, PJ, SP, Tangerang 16 1e-4 20
@@ C.3. Experiment and Implementation Details
STHGCN Beijing, Melbourne, Palembang, Shanghai, Sydney, Tokyo 16 1e-4 20
Bandung, Istanbul, Jakarta, KC, Moscow, NY, PJ, SP, Tangerang 64 1e-4 20
@@
All modified code implementations are available as submodules in our main dataset repository. Experiments were conducted using NVIDIA L4, L40S, and H100 GPUs.
C.4. Supplementary Results
@@
We report the full results of our supervised POI recommendation baselines in Table 12, 13 and 14, using three evaluation metrics: Acc@1, Acc@5, and NDCG@5.
@@
Table 12. Performance of supervised POI recommendation baselines across 5 cities: Bandung, Beijing, Istanbul, Jakarta, Kuwait City. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model Bandung Beijing Istanbul Jakarta Kuwait City
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.048 0.118 0.083 0.000 0.021 0.009 0.026 0.074 0.050 0.029 0.085 0.058 0.021 0.089 0.054
RNN 0.062 0.135 0.099 0.085 0.183 0.134 0.077 0.178 0.130 0.049 0.115 0.083 0.087 0.203 0.146
LSTPM 0.110 0.241 0.179 0.127 0.211 0.169 0.142 0.286 0.217 0.099 0.210 0.157 0.180 0.362 0.275
DeepMove 0.107 0.232 0.172 0.106 0.261 0.190 0.150 0.298 0.228 0.103 0.212 0.160 0.179 0.360 0.274
Flashback 0.103 0.202 0.155 0.278 0.435 0.359 0.137 0.257 0.200 0.113 0.214 0.166 0.150 0.298 0.227
Graph-Flashback 0.091 0.197 0.146 0.304 0.426 0.373 0.137 0.261 0.202 0.072 0.167 0.121 0.115 0.245 0.181
LoTNext 0.076 0.169 0.125 0.243 0.391 0.320 0.105 0.227 0.168 0.044 0.110 0.078 0.132 0.280 0.209
GETNext 0.179 0.306 0.247 0.433 0.527 0.486 0.146 0.268 0.210 0.155 0.257 0.209 0.175 0.322 0.251
STHGCN 0.219 0.375 0.302 0.453 0.640 0.552 0.241 0.385 0.318 0.197 0.334 0.270 0.225 0.394 0.314
UniMove 0.007 0.060 0.033 0.036 0.205 0.128 0.015 0.061 0.038 0.004 0.036 0.020 0.023 0.120 0.073
@@
Table 13. Performance of supervised POI recommendation baselines across 5 cities: Melbourne, Moscow, New York, Palembang, Petaling Jaya. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model Melbourne Moscow New York Palembang Petaling Jaya
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.062 0.147 0.107 0.059 0.129 0.094 0.032 0.090 0.061 0.102 0.169 0.136 0.026 0.084 0.057
RNN 0.059 0.105 0.083 0.075 0.164 0.122 0.061 0.119 0.092 0.049 0.121 0.085 0.064 0.148 0.107
LSTPM 0.091 0.204 0.150 0.151 0.300 0.229 0.099 0.206 0.155 0.114 0.230 0.175 0.099 0.222 0.163
DeepMove 0.083 0.179 0.134 0.143 0.283 0.217 0.097 0.195 0.149 0.084 0.191 0.139 0.112 0.234 0.175
Flashback 0.079 0.146 0.115 0.121 0.234 0.180 0.086 0.156 0.123 0.135 0.256 0.198 0.072 0.162 0.119
Graph-Flashback 0.093 0.188 0.143 0.125 0.243 0.188 0.059 0.125 0.093 0.148 0.264 0.208 0.072 0.164 0.119
LoTNext 0.085 0.180 0.133 0.089 0.196 0.145 0.061 0.133 0.099 0.140 0.271 0.208 0.055 0.135 0.097
GETNext 0.100 0.250 0.179 0.175 0.335 0.260 0.134 0.263 0.202 0.158 0.313 0.239 0.139 0.254 0.200
STHGCN 0.168 0.318 0.247 0.223 0.382 0.308 0.146 0.259 0.207 0.246 0.427 0.341 0.174 0.301 0.241
UniMove 0.008 0.066 0.037 0.009 0.051 0.030 0.004 0.028 0.016 0.009 0.060 0.035 0.008 0.058 0.034
Table 14. Performance of supervised POI recommendation baselines across 5 cities: São Paulo, Shanghai, Sydney, Tangerang, Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model São Paulo Shanghai Sydney Tangerang Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.030 0.079 0.055 0.084 0.154 0.120 0.075 0.180 0.131 0.104 0.220 0.166 0.176 0.291 0.239
RNN 0.097 0.191 0.147 0.055 0.120 0.090 0.080 0.164 0.125 0.087 0.179 0.135 0.133 0.254 0.197
LSTPM 0.158 0.319 0.243 0.099 0.195 0.149 0.141 0.265 0.206 0.154 0.309 0.237 0.225 0.394 0.315
DeepMove 0.160 0.310 0.240 0.085 0.168 0.128 0.129 0.240 0.188 0.145 0.285 0.219 0.201 0.362 0.288
Flashback 0.159 0.280 0.222 0.096 0.162 0.132 0.127 0.216 0.174 0.185 0.330 0.262 0.170 0.298 0.237
Graph-Flashback 0.096 0.211 0.156 0.111 0.187 0.151 0.135 0.234 0.188 0.200 0.373 0.292 0.180 0.329 0.259
LoTNext 0.105 0.222 0.166 0.093 0.179 0.139 0.123 0.225 0.176 0.126 0.249 0.190 0.191 0.315 0.258
GETNext 0.202 0.360 0.286 0.115 0.230 0.177 0.181 0.347 0.266 0.224 0.372 0.302 0.180 0.361 0.275
STHGCN 0.250 0.425 0.344 0.193 0.329 0.264 0.227 0.378 0.307 0.293 0.492 0.400 0.250 0.432 0.350
UniMove 0.002 0.018 0.009 0.000 0.055 0.029 0.015 0.102 0.059 0.001 0.055 0.029 0.032 0.109 0.072
@@ Appendix D Analyzing Urban Features and POI Recommendation Performance
POI trajectory studies that evaluate models across multiple city-level datasets often include analyses to assess how well their methods generalize across different urban contexts. It is well understood that POI recommendation accuracy metrics (e.g., Acc@k, NDCG@k) can vary substantially between cities and can be interpreted as a proxy for how easy or difficult a city is to model. The assumption is that higher performance reflects more predictable or structured mobility patterns. This viewpoint is consistent with prior work highlighting the role of cultural and urban-specific factors in shaping mobility behaviors (Yang et al., 2015; Sun et al., 2024). For example, Gowalla-CA (Cho et al., 2011; Yuan et al., 2013) often yields lower accuracy compared to FSQ-NYC and FSQ-TKY (Yang et al., 2014), suggesting that some cities may be inherently harder to model.
Several studies have proposed hypotheses connecting specific urban features to modeling difficulty. Yang et al. (2022c) hypothesized that cities with fewer check-ins and higher spatial sparsity of POIs are harder to model. Yan et al. (2023) suggested that a larger number of user trajectories improves predictive accuracy by providing richer collaborative signals, whose architecture is designed to leverage. Li et al. (2024) proposed that cities with a greater variety of POI categories are easier to model due to LLMs’ contextual reasoning capabilities, whereas cities covering a broader geographic area tend to be more difficult to model. In the zero-shot POI recommendation setting, Feng et al. (2025) reported two key findings: (1) geospatial biases inherent in LLMs can hinder prediction quality across cities, and (2) LLMs are influenced by city-specific mobility patterns.
Building on these insights, we used Massive-STEPS to explore how urban features affect POI recommendations. Its diverse set of 15 cities allows for a comprehensive analysis across different cultural and urban contexts. We analyzed the correlation between urban features and model accuracy, and based on the results, proposed a new hypothesis that contrasts previous findings in the literature.
Prior studies (Yang et al., 2022c; Yan et al., 2023; Li et al., 2024) have suggested several features as potential explanatory variables, including:
@@
Moreover, previous studies have primarily focused on only three datasets: FSQ-NYC, FSQ-TKY, and Gowalla-CA. In contrast, Massive-STEPS provides broader coverage across 15 cities, enabling a more comprehensive and robust analysis. To examine the relationship between urban features and model performance, we averaged the three evaluation metrics across nine supervised baselines for each city and computed the Spearman correlation with each candidate feature. Fig. 11 presents the correlations between all nine features and the average performance metric.
Among all features, category entropy shows the strongest correlation with model performance, with a Spearman correlation of r=0.646r=-0.646 ($p<0.05$). This suggests that cities with more evenly distributed POI categories tend to yield lower prediction accuracy. Intuitively, when no single category dominates (a city has roughly equal proportions of restaurants, cafes, homes, and other POIs), it becomes more difficult for models to predict a user’s next destination. In these cases, user behavior is more varied and less predictable. On the other hand, cities with more skewed category distributions (e.g., mostly food places or mostly residential areas) tend to have more consistent patterns of movement, making them easier for models to learn and predict. Interestingly, our finding contradicts the hypothesis proposed by LLM4POI (Li et al., 2024), which suggests that FSQ-NYC is easier to model than Gowalla-CA due to the former’s vast number of POI categories, which were supposed to provide richer contextual signals for the model. Our results indicate that it is not the number of categories that matters, but rather how these categories are distributed.
Figure 11. Spearman correlation between nine candidate urban features and the mean score of POI recommendation models across 15 cities.
Spearman correlation between nine candidate urban features and the mean score of POI recommendation models across 15 cities.
@@ E.1. Problem Formulation
E.2. Methods
We evaluated three LLM-based prompting methods:
  • LLM-Mob (Wang et al., 2023c): One of the earliest methods to use LLMs for next POI prediction, prompting LLMs with both historical and current (contextual) trajectories.
  • LLM-ZS (Beneduce et al., 2024): A simplified version of LLM-Mob that retains the use of historical and contextual trajectories but simplifies its prompt design.
  • LLM-Move (Feng et al., 2024a): Extends previous prompting methods by introducing a RAG-like approach, retrieving nearby POIs as candidates, and ranking them by geographic distance to the user’s most recent visit.
E.3. Experiment and Implementation Details